# An analogue of Reed’s conjecture for digraphs

### v2

@@ 6 Proof of the main result
$\Delta_{\max}\geq\tilde{\Delta}\geq\max(\Delta_{1},\Delta_{2},\Delta(a))$.
Proof of claim. It is clear from the definitions that $\Delta_{\max}\geq\tilde{\Delta}$. Assume for a contradiction that $\tilde{\Delta}<\max(\Delta_{1},\Delta_{2},\Delta(a))$. Since $\varepsilon\leq\min\left(\frac{1}{\Delta_{1}},\frac{1}{\Delta_{2}},\frac{1}{\Delta(a)}\right)$ we have $\varepsilon<\frac{1}{\tilde{\Delta}}$, which implies:
$k=\left\lceil\tilde{\Delta}+1+\varepsilon(\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}-\tilde{\Delta}-1)\right\rceil\geq\left\lceil\tilde{\Delta}+\frac{1}{\tilde{\Delta}}(\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}-1)\right\rceil.$
Using the facts that $\vec{\chi}(D)>k$ and that $\vec{\chi}(D)\leq\left\lfloor\tilde{\Delta}+1\right\rfloor$, we deduce
@@
This is a contradiction unless ${\omega}=1$ and $\tilde{\Delta}$ is an integer. Henceforth we assume that ${\omega}=1$, i.e. $D$ is an oriented graph, and $\tilde{\Delta}\in\mathbb{N}$. In particular, we have $k\geq\tilde{\Delta}$. As $D$ is an oriented graph, by Theorem 6.1 we must have $\Delta_{\min}(D)\leq 1$, for otherwise $\vec{\chi}(D)\leq\Delta_{\min}(D)\leq\tilde{\Delta}\leq k$, a contradiction. If $\tilde{\Delta}=1$ then $k=\lceil 2-\varepsilon\rceil=2$. If $\tilde{\Delta}=0$ then $k=1$. Both cases yield a contradiction, which concludes the proof of the claim. $\lozenge$
@@
Case 1
There exists a vertex $v$ such that $\max(m^{-}(v),m^{+}(v))\geq(1-a)\Delta(\Delta-1)$.
By Claim 6.3 we have ${\omega}\leq\frac{2}{3}(\Delta_{\max}+1)$. As $\varepsilon_{0}=\frac{1}{6}-4\sqrt{a}$, by Theorem 1.9 we have
$\vec{\chi}(D)\leq\max(\vec{\chi}(D-v),(1-\varepsilon_{0})(\Delta_{\max}+1)).$
Recall that $\vec{\chi}(D-v)\leq k$ as $D$ is a minimum counterexample. Since $\vec{\chi}(D)>k$, we thus necessarily have
$\vec{\chi}(D)\leq(1-\varepsilon_{0})(\Delta_{\max}+1)\leq(1-\varepsilon_{0})(1+\gamma_{\varepsilon})(\tilde{\Delta}+1)\leq(1-\varepsilon)(\tilde{\Delta}+1),$
where in the second inequality we used Claim 6.2, and in the last inequality we used $\varepsilon\leq\varepsilon_{0}-\gamma_{\varepsilon}$, which holds because $\gamma_{\varepsilon}\leq 2\varepsilon$ and $\varepsilon\leq\frac{1}{3}\varepsilon_{0}$. This yields the contradiction as $(1-\varepsilon)(\tilde{\Delta}+1)\leq k$.
Case 2
For every vertex $v\in V(D)$ we have $\max(m^{-}(v),m^{+}(v))\leq(1-a)\Delta(\Delta-1)$.
A fortiori every vertex satisfies $\min(m^{-}(v),m^{+}(v))\leq(1-a)\Delta(\Delta-1)$. By Claim 6.4 we know that $\Delta_{\max}\geq\max(\Delta_{1},\Delta_{2})$. Hence, by definition of $\Delta_{2}$, we have $a>\frac{\log^{3}\Delta}{\Delta-1}$. We can thus apply Corollary 3.2, which together with Claim 6.2 gives
$\vec{\chi}(D)\leq\left(1-\frac{a}{5e^{7}}\right)(\Delta_{\max}+1)\leq\left(1-\frac{a}{5e^{7}}\right)(1+\gamma_{\varepsilon})(\tilde{\Delta}+1)\leq\left(1-\varepsilon\right)(\tilde{\Delta}+1).$
where in the last inequality we used $\varepsilon\leq\frac{a}{5e^{7}}-\gamma_{\varepsilon}$, which holds because $\gamma_{\varepsilon}\leq 2\varepsilon$ and $\varepsilon\leq\frac{a}{15e^{7}}$. This contradicts $\vec{\chi}(D)>k$ and concludes the proof. ∎
@@ 7 Further research: analogues for $\Delta_{\min}$
$\vec{\chi}(D)>\lceil(1-\varepsilon)(\Delta_{\min}(D)+1)+\varepsilon^{\prime}\Delta_{\min}(D)\rceil.$
This is because the $k$-Dicolourability problem is NP-complete for every fixed $k\geq 2$, even when restricted to digraphs $D$ satisfying $\Delta_{\min}(D)=k$ and ${\omega}(D)=\left\lceil\frac{k+1}{2}\right\rceil$, as shown in [29]. In particular, the reduction shows that there exist such digraphs satisfying $\vec{\chi}(D)=k+1$. Hence, if we choose $k$ to be at least $\frac{2-2\varepsilon+2\varepsilon^{\prime}}{2\varepsilon-\varepsilon^{\prime}}$, we deduce the existence of digraphs $D$ with $\Delta_{\min}(D)=k$ and $\vec{\chi}(D)=k+1$, while $\lceil(1-\varepsilon)(k+1)+\varepsilon^{\prime}k\rceil\leq k$.
@@
$\vec{\chi}(D)\leq\lceil(1-\varepsilon)\Delta_{\min}(D)+\varepsilon\vec{\omega}(D)\rceil.$
@@
$\vec{\chi}(D)\leq\lceil(1-\varepsilon)\Delta_{\min}(D)+2\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\rceil.$
We are inclined to believe that these two conjectures are true even for $\varepsilon=\frac{1}{2}$. One direction for proving Conjecture 7.1 would be to prove an analogue of Theorem 1.6 using the maximum out-degree $\Delta^{+}$ instead of $\tilde{\Delta}$, that is to prove the existence of $\varepsilon>0$ such that every digraph $D$ satisfies
$\vec{\chi}(D)\leq\lceil(1-\varepsilon)(\Delta^{+}(D)+1)+\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\rceil,$
@@
For every digraph $D$, there exists a digraph $H$ such that $\vec{\chi}(D)\leq\vec{\chi}(H)$, $\Delta^{+}(H)\leq\Delta_{\min}(D)$ and ${\omega}(H)\leq\vec{\omega}(D)$.
@@
Let us fix a digraph $D$, and let $(X,Y)$ be the partition of $V(D)$ where $X=\{x\in V(D)\mid d^{+}_{D}(v)\leq\Delta_{\min}(D)\}$ and $Y=V(D)\setminus X$. In particular, every vertex $y\in Y$ satisfies $d^{-}(y)\leq\Delta_{\min}(D)$.
Let $D^{\prime}$ be the digraph obtained from $D$ by removing every arc from $Y$ to $X$ and replacing every arc from $X$ to $Y$ by a digon. Let $H$ be the digraph obtained from $D^{\prime}$ by reversing every arc of $D^{\prime}[Y]$. It is straightforward to check that, by construction, $\Delta^{+}(H)\leq\Delta_{\min}(D^{\prime})$ and ${\omega}(H)\leq\vec{\omega}(D)$.
We now justify that $\vec{\chi}(H)\geq\vec{\chi}(D)$. Assume that this is not the case, and let $\phi$ be a $(\vec{\chi}(D)-1)$-dicolouring of $H$. By definition, $D$ coloured with $\phi$ must contain a monochromatic directed cycle $C=v_{1},\dots,v_{\ell},v_{1}$. Observe that $V(C)\nsubseteq X$, otherwise $C$ is a directed cycle of $H$. In a similar way, $V(C)\nsubseteq Y$, for otherwise $v_{1},v_{\ell},\dots,v_{2},v_{1}$ is a directed cycle of $H$. Hence $C$ contains an arc $xy$ from $X$ to $Y$, but $\{xy,yx\}$ is a digon of $H$, a contradiction. ∎
@@
Problem 7.4.
Show the existence of $\varepsilon>0$ such that every oriented graph $D$ with sufficiently large maximum out-degree $\Delta^{+}(D)$ satisfies $\vec{\chi}(D)\leq(1-\varepsilon)\Delta^{+}(D)$.
@@ Acknowledgments
The first author is supported by JSPS KAKENHI JP20A402 and 22H05001 and by JST ASPIRE JPMJAP2302. The second author is supported by the research grants DIGRAPHS ANR-19-CE48-0013 and ANR-17-EURE-0004.

### v3

@@ 7 Proof of the main result
$\Delta_{\max}\geqslant\tilde{\Delta}\geqslant\max(\Delta_{1},\Delta_{2},\Delta(a))$.
Proof of claim. It is clear from the definitions that $\Delta_{\max}\geqslant\tilde{\Delta}$. Assume for a contradiction that $\tilde{\Delta}<\max(\Delta_{1},\Delta_{2},\Delta(a))$. Since $\varepsilon\leqslant\min\left(\frac{1}{\Delta_{1}},\frac{1}{\Delta_{2}},\frac{1}{\Delta(a)}\right)$ we have $\varepsilon<\frac{1}{\tilde{\Delta}}$, which implies:
$k=\left\lceil\tilde{\Delta}+1+\varepsilon(\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}-\tilde{\Delta}-1)\right\rceil\geqslant\left\lceil\tilde{\Delta}+\frac{1}{\tilde{\Delta}}(\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}-1)\right\rceil,$
where in the inequality above we used that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}\leqslant\tilde{\Delta}+1$. This holds as any vertex $v$ in a maximum biclique has in-degree and out-degree at least $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}-1$, hence implying that
$\tilde{\Delta}+1\geqslant\sqrt{(\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}-1)^{2}}+1=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}.$
Using the facts that $\vec{\chi}(D)>k$ and that $\vec{\chi}(D)\leqslant\left\lfloor\tilde{\Delta}+1\right\rfloor$, we deduce
@@
This is a contradiction unless $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}=1$ and $\tilde{\Delta}$ is an integer. Henceforth we assume that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}=1$, i.e., $D$ is an oriented graph, and $\tilde{\Delta}\in\mathbb{N}$. In particular, we have $k\geqslant\tilde{\Delta}$. As $D$ is an oriented graph, by Theorem 7.1 we must have $\Delta_{\min}(D)\leqslant 1$, for otherwise $\vec{\chi}(D)\leqslant\Delta_{\min}(D)\leqslant\tilde{\Delta}\leqslant k$, a contradiction. If $\tilde{\Delta}=1$ then $k=\lceil 2-\varepsilon\rceil=2$. If $\tilde{\Delta}=0$ then $k=1$. Both cases yield a contradiction, which concludes the proof of the claim. $\lozenge$
@@
Case 1:
There exists a vertex $v$ such that $\max(m^{-}(v),m^{+}(v))\geqslant(1-a)\Delta(\Delta-1)$.
By Claim 7.3 we have $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}\leqslant\frac{2}{3}(\Delta_{\max}+1)$. As $\varepsilon_{0}=\frac{1}{6}-4\sqrt{a}$, by Theorem 1.10 we have
$\vec{\chi}(D)\leqslant\max(\vec{\chi}(D-v),(1-\varepsilon_{0})(\Delta_{\max}+1)).$
Recall that $\vec{\chi}(D-v)\leqslant k$ as $D$ is a minimum counterexample. Since $\vec{\chi}(D)>k$, we thus necessarily have
$\vec{\chi}(D)\leqslant(1-\varepsilon_{0})(\Delta_{\max}+1)\leqslant(1-\varepsilon_{0})(1+\gamma_{\varepsilon})(\tilde{\Delta}+1)\leqslant(1-\varepsilon)(\tilde{\Delta}+1),$
where in the second inequality we used Claim 7.2, and in the last inequality we used $\varepsilon\leqslant\varepsilon_{0}-\gamma_{\varepsilon}$, which holds because $\gamma_{\varepsilon}\leqslant 2\varepsilon$ and $\varepsilon\leqslant\frac{1}{3}\varepsilon_{0}$. This yields the contradiction as $(1-\varepsilon)(\tilde{\Delta}+1)\leqslant k$.
Case 2:
For every vertex $v\in V(D)$ we have $\max(m^{-}(v),m^{+}(v))\leqslant(1-a)\Delta(\Delta-1)$.
A fortiori every vertex satisfies $\min(m^{-}(v),m^{+}(v))\leqslant(1-a)\Delta(\Delta-1)$. By Claim 7.4 we know that $\Delta_{\max}\geqslant\max(\Delta_{1},\Delta_{2})$. Hence, by definition of $\Delta_{2}$, we have $a>\frac{\log^{3}\Delta}{\Delta-1}$. We can thus apply Corollary 4.2, which together with Claim 7.2 gives
$\vec{\chi}(D)\leqslant\left(1-\frac{a}{5e^{7}}\right)(\Delta_{\max}+1)\leqslant\left(1-\frac{a}{5e^{7}}\right)(1+\gamma_{\varepsilon})(\tilde{\Delta}+1)\leqslant\left(1-\varepsilon\right)(\tilde{\Delta}+1).$
where in the last inequality we used $\varepsilon\leqslant\frac{a}{5e^{7}}-\gamma_{\varepsilon}$, which holds because $\gamma_{\varepsilon}\leqslant 2\varepsilon$ and $\varepsilon\leqslant\frac{a}{15e^{7}}$. This contradicts $\vec{\chi}(D)>k$ and concludes the proof. ∎
@@ 8 Further research: analogues for $\Delta_{\min}$
$\vec{\chi}(D)>\lceil(1-\varepsilon)(\Delta_{\min}(D)+1)+\varepsilon^{\prime}\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\rceil.$
This is because the $k$-Dicolourability problem is NP-complete for every fixed $k\geqslant 2$, even when restricted to digraphs $D$ satisfying $\Delta_{\min}(D)=k$ and $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)=\left\lceil\frac{k+1}{2}\right\rceil$, as shown in [32]. In particular, the reduction shows that there exist such digraphs satisfying $\vec{\chi}(D)=k+1$. Hence, if we choose $k$ to be at least $\frac{2-2\varepsilon+2\varepsilon^{\prime}}{2\varepsilon-\varepsilon^{\prime}}$, we deduce the existence of digraphs $D$ with $\Delta_{\min}(D)=k$ and $\vec{\chi}(D)=k+1$, while $\lceil(1-\varepsilon)(k+1)+\varepsilon^{\prime}k\rceil\leqslant k$.
@@
$\vec{\chi}(D)\leqslant\lceil(1-\varepsilon)\Delta_{\min}(D)+\varepsilon\vec{\omega}(D)\rceil.$
@@
$\vec{\chi}(D)\leqslant\lceil(1-\varepsilon)\Delta_{\min}(D)+2\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\rceil.$
We are inclined to believe that these two conjectures are true even for $\varepsilon=\frac{1}{2}$. One direction toward proving Conjecture 8.1 would be to prove an analogue of Theorem 1.7 using the maximum out-degree $\Delta^{+}$ instead of $\tilde{\Delta}$, that is to prove the existence of $\varepsilon>0$ such that every digraph $D$ satisfies
$\vec{\chi}(D)\leqslant\lceil(1-\varepsilon)(\Delta^{+}(D)+1)+\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\rceil,$
@@
For every digraph $D$, there exists a digraph $H$ such that $\vec{\chi}(D)\leqslant\vec{\chi}(H)$, $\Delta^{+}(H)\leqslant\Delta_{\min}(D)$ and $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(H)\leqslant\vec{\omega}(D)$.
@@
Let us fix a digraph $D$, and let $(X,Y)$ be the partition of $V(D)$ where $X=\{x\in V(D)\mid d^{+}_{D}(v)\leqslant\Delta_{\min}(D)\}$ and $Y=V(D)\setminus X$. In particular, every vertex $y\in Y$ satisfies $d^{-}(y)\leqslant\Delta_{\min}(D)$.
Let $D^{\prime}$ be the digraph obtained from $D$ by removing every arc from $Y$ to $X$ and replacing every arc from $X$ to $Y$ by a digon. Let $H$ be the digraph obtained from $D^{\prime}$ by reversing every arc of $D^{\prime}[Y]$. It is straightforward to check that, by construction, $\Delta^{+}(H)\leqslant\Delta_{\min}(D^{\prime})$ and $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(H)\leqslant\vec{\omega}(D)$.
We now justify that $\vec{\chi}(H)\geqslant\vec{\chi}(D)$. Assume that this is not the case, and let $\phi$ be a $(\vec{\chi}(D)-1)$-dicolouring of $H$. By definition, $D$ coloured with $\phi$ must contain a monochromatic directed cycle $C=v_{1},\dots,v_{\ell},v_{1}$. Observe that $V(C)\nsubseteq X$, otherwise $C$ is a directed cycle of $H$. In a similar way, $V(C)\nsubseteq Y$, for otherwise $v_{1},v_{\ell},\dots,v_{2},v_{1}$ is a directed cycle of $H$. Hence $C$ contains an arc $xy$ from $X$ to $Y$, but $\{xy,yx\}$ is a digon of $H$, a contradiction. ∎
@@
Problem 8.4.
Show the existence of $\varepsilon>0$ such that every oriented graph $D$ with sufficiently large maximum out-degree $\Delta^{+}(D)$ satisfies $\vec{\chi}(D)\leqslant(1-\varepsilon)\Delta^{+}(D)$.
@@ Acknowledgments
The first and second authors are supported by JSPS KAKENHI JP20A402 and 22H05001 and by JST ASPIRE JPMJAP2302. The second author has been supported by the research grants DIGRAPHS ANR-19-CE48-0013 and ANR-17-EURE-0004.